OpenStudy (anonymous):

Let f(x) = 2x+3 and g(x) = x2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x)) Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2.

OpenStudy (anonymous):

In English, f(g(x)) means take the expression for g(x) (in your example that would be "x2") an replace every "x" in the f(x) with your g(x) expression.

OpenStudy (anonymous):

Take f(x) = 2x + 3. Now replace the x with x2 and simplify. What do you get?

OpenStudy (anonymous):

Assuming g(x) is x^2... 1. f(g(x)) = 2x^2 + 3 2. (2x + 3)^2 = 4x^2 + 12x + 9 3. The difference is which term is substituted for the variable x. f(g(x)) means substitute the entire expression g(x) for every variable x in the expression f(x), and vice-versa..

OpenStudy (anonymous):

Yes its ^2 sorry about that

OpenStudy (anonymous):

Nothign else could be done to part 1 right?

OpenStudy (anonymous):

In part 1, just substitute x^2 for x in f(x) ... 2*(x^2) + 3 = 2x^2 + 3

OpenStudy (anonymous):

thank you